OpenStudy (callisto):

If f(x) is divisible by 2x+3, then f(x+1) is divisible by?

OpenStudy (anonymous):

2x + 5? hehe just guessing

OpenStudy (callisto):

i forgot, I just suddenly remember this type of question and I cannot solve it..

OpenStudy (anonymous):

1

OpenStudy (anonymous):

2(x+1) + 3 ?

OpenStudy (anonymous):

From the given information, f(x) = (2x+3) Q(x) so, f(x+1) = (2x+5) Q(x+1)

OpenStudy (anonymous):

So, Ishaan is right.

OpenStudy (apoorvk):

Yeah Ishaan, I would have guessed the same.

OpenStudy (callisto):

Alright , i don't understand. What about if f(x) is divisible by x-1, which of the following must be a factor of f(2x+1) ? A. x B . x-3 C. 2x-1 D. 2x+1 This time.. I've got the answer for this one , but i don't know how to solve it

OpenStudy (anonymous):

A

OpenStudy (anonymous):

same method.

OpenStudy (anonymous):

Let \( f (x) = (x - 1) Q (x)\), then \[ f (2x + 1) = (2x + 1 - 1) Q (2x + 1)= 2x Q (2x + 1) \] Hence x is a factor of f (2x + 1). (Ans: A)

OpenStudy (callisto):

Oh this time i understand, thanks!!!

OpenStudy (anonymous):

I am glad you did.

OpenStudy (anonymous):

Time for the problem of the day guys! ;)

OpenStudy (callisto):

right, for the first one i don't understand i thought f(x) = (2x+3) Q(x) f(x+1) = (2x+3 +1) Q(x+1) = 2(x+1) Q(x+1) ?

OpenStudy (mani_jha):

When you find f(x+1), you just substitute x+1 instead of x. No other changes in that. Just like to find f(2) you substitute 2 instead of x. So f(x+1)={2(x+1)+3)Q(x+1) ={2x+2+3}Q(x+1) =(2x+5)Q(x+1) Ok?

OpenStudy (callisto):

Yup, thanks a ton!